# Affordance-Aware Object Insertion via Mask-Aware Dual Diffusion

### v1

@@ A.3 Data Augmentation
Table 5: Reserved percentage for foreground quality control filters. We combine different rule-based and learning-based conditions. Through this process, foreground objects with high quality are reserved.
Figure 9: FID-CLIP score curve on $128\times 128$ resolution with different guidance scale $[1.0,3.0,4.0,5.0,6.0,7.0]$.
@@
Figure 10: 10(a) shows the candidate foreground samples in the pipeline. The upper row shows four low-quality samples. The lower row shows the samples after data quality control. 10(b) shows the word cloud of foreground categories in the SAM-FB dataset.
@@
Baseline 25.89 26.21 26.37 27.35 26.46 89.12 89.50 79.92 79.31 84.46
+Classifier-Free 21.93 22.03 22.31 22.74 22.25 91.13 90.95 85.49 85.26 88.21
+Dual Diffusion 21.75 21.81 21.90 22.39 21.96 91.57 91.05 88.25 88.34 89.80
+Expertise branch 21.55 21.66 21.76 22.24 21.80 91.68 90.96 89.61 88.30 90.14
@@
A.4 Classifier Free Guidance
An essential mechanism we employ to support the Null position prompt is Classifier-Free Guidance (CFG). This approach not only enhances the performance of a well-trained Stable Diffusion model but also simplifies handling Null prompts by converting them into Classifier-Free cases. It does this by controlling additional guidance signals each time we sample $\mathbf{z}_{t-1}$ from $\mathbf{z}_{t}$ and $\hat{\boldsymbol{\epsilon}}_{t}$.
Ho & Salimans[16] figured out that it is feasible to use the same model to generate an inherent Classifier-Free Guidance signal $p(\mathbf{z}_{t}|\varnothing)$ as long as we drop certain conditions during the training procedure. Since we have three different conditions $\mathbf{f}$, $\mathbf{b}$, and $\mathbf{p}$, we drop all the conditions with a 0.1 probability and only drop $\mathbf{p}$ with a 0.1 probability. During the inference steps, we can use the CFG to guide the sampling schedule as follows:
$\hat{\boldsymbol{\epsilon}}_{t}=(1+s)\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t};\mathbf{b},\mathbf{f},\mathbf{p},t)-s\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t};\varnothing,t)$ (8)
where $s$ is the classifier-free guidance scale.
We test the FID and CLIP scores for different prompts with different guidance scales $s$. Figure  9 shows the FID-CLIP curves using different guidance scales ranging from $[1.0,3.0,4.0,5.0,6.0,7.0]$. We set the guidance scale to 4.0 as it usually gives good results across all different prompts.
@@ B.2 Human Evaluation
We conducted our user survey by asking 10 users to rank outputs from different models in 10 affordance insertion settings, therefore gaining 100 data points.
@@ C.1 Details Maintaining
Using DINOv2 features, the model preserves the appearance more effectively than other image editing models, especially for objects with detailed textures. Figure 11 compares our model with other image editing models. Our model demonstrates the ability to retain the details of an object’s appearance, even the texture of the object.
@@ C.2 SAM-FB Test Split
Figure 12 shows a visualization result with different baselines on the SAM-FB test split. Our model generated more authentic results compared to other baseline models.
Figure 12: Samples on SAM-FB test split. Our model inserted the bag with an authentic appearance.
@@ C.3 Detailed Ablations
Figure 13: Example of in-the-wild insertion results with details. Our model could keep the appearance better and adjust the foreground’s properties better compared with SD [39], GLI-GEN [29] and PBE [47] on common objects. In the last row, the model generated reasonable insertion when provided with ambiguous prompts.
C.4 In-the-wild Generalization
We compare our methods with other baseline models on in-the-wild images, and Figure 13 shows the visualization results for comparison. In the first five rows, we performed the affordance insertion task providing a bounding box on both common objects like apples and uncommon objects like cruta. Our methods generated the images with the highest quality, authentic to the reference foreground with proper lighting. In the last row, we show more affordance insertion results when providing ambiguous prompts. The model will automatically find the proper affordance relationship and adjust the location and view of the foreground object. It’s notable that in the first case when inserting the cruise, it is actually the view from the back of the reference image.
C.5 Video Demo
Please refer to the video demo on the project website https://kakituken.github.io/affordance-any.github.io/.
C.6 Failure Cases
Figure 14 shows some failure cases when using the model to perform affordance insertion. Generally, when prompted null position, it requires the model to search for a possible position to insert the object. However, if there are already similar objects in the scene e.g., a traffic light in the first row, it is easy to mislead the model and end up inserting nothing. When the background is too complex and the foreground object is too small, such as a vegetable in a supermarket shown in the second row, it is also difficult for the model to insert the object correctly.

### v2

@@ 3.2 Dataset Construction Pipeline
Table 2: Reserved percentages for foreground quality control filters. We apply a combination of rule-based and learning-based conditions to ensure high-quality foreground objects are retained through a rigorous filtering process.
Foreground Filtering. Although NMS removes some low-quality masks, SAM segmentation can still produce unintended elements due to challenges in controlling granularity and semantic consistency. The resulting masks often suffer from three main issues: Incompleteness: Some objects are partially occluded, resulting in fragmented masks that only capture small portions of the original object (e.g., half of a tree); Background Masks: Some masks mistakenly segment background elements (e.g., sky, grassland); Undersized / Oversized Objects: Very small objects often have blurry details, reducing dataset quality. To enhance foreground quality, we employ a hybrid quality control (QC) approach combining rule-based and learning-based filtering.
For rule-based filtering, we remove low-quality masks with four constraints: relative size (too small/large objects), aspect ratio (overly thin objects), component count (disjointed objects), and color standard deviation (pure-color backgrounds). While rule-based filtering effectively removes outliers, ensuring object completeness is challenging using predefined rules alone. To address this, we manually annotate 2,000 foreground images with binary labels (good vs. bad) and fine-tune a ResNet-50 [16] classifier for binary classification. The trained classifier assigns a quality score to each foreground object, further refining the dataset. Table 2 demonstrates the specific threshold for each filter operation of our data quality control stage. With these filter operations, only 0.25% of the masks are left, which ensures the quality of the obtained dataset.
Background Inpainting. High-quality foregrounds are then cropped from the original images. Instead of covering the removed regions with opaque masks, which act as strong artificial cues for object position and size, we employ LAMA [45] for background inpainting. LAMA is a simple yet effective method for object removal and hole-filling. Following GroundingSAM [32, 41], we expand the object mask boundary slightly to prevent residual foreground content from remaining in the background. By eliminating such visual cues, our dataset encourages models to infer context-aware object placement based on scene affordances.
Background Inpainting. Despite careful inpainting, some regions inevitably contain inpainting artifacts. To assess the quality of inpainted backgrounds, we compute the Structural Similarity Index (SSIM) between the inpainted background and the original image. We discard any inpainted background with an SSIM score below a certain threshold to maintain high visual consistency. We empirically set the SSIM threshold to 0.8 based on preliminary experiments assessing inpainting realism. If a background fails the SSIM filter, we discard the entire foreground-background pair, even if the foreground quality is high. This ensures that only high-quality samples are included in the dataset.
To evaluate quality improvements, we compute the Inception Score (IS) for foregrounds and Fréchet Inception Distance (FID) for backgrounds, comparing values before and after quality control (QC). After QC, foreground IS improves from 9.55 to 16.68, indicating significantly enhanced foreground clarity, while background FID decreases from 9.23 to 8.46, reflecting better distribution consistency, similar to inpainting tasks. Since FID requires a reference distribution, it is less meaningful for foregrounds, whereas IS is not a reliable metric for complex backgrounds. To justify the diversity of our proposed dataset, we use RAM [53] to recognize the categories of foregrounds we keep. We have identified 3439 different categories, which is a vast improvement over previous datasets. There are no manual annotations required in the pipeline, and users can easily scale up the dataset further with any new source images easily.
4 Method
4.1 Mask-Aware Dual Diffusion Model (MADD)
Preliminaries. Given a foreground object $\boldsymbol{f}$, a background scene $\boldsymbol{b}$, and a position prompt $\boldsymbol{p}$, the goal is to synthesize an image $\boldsymbol{x}$ that integrates the object into the scene, while adhering to affordances and aligning with the position prompt.
To develop an effective diffusion model, we first extract image features $\boldsymbol{z}=\mathcal{E}(x)$ using a pre-trained encoder. Then we predict the object insertion mask $\boldsymbol{m}$ capturing object’s location and size, which can be directly derived from the difference between $\boldsymbol{x}$ and $\boldsymbol{b}$ in the training data. Our MADD model $\mathcal{G}$ builds upon the latent structural diffusion framework [33], simultaneously denoising $\boldsymbol{\hat{z}}$ and $\boldsymbol{\hat{m}}$ (Figure 2):
$\hat{\boldsymbol{z}},\hat{\boldsymbol{m}}=\mathcal{G}(\boldsymbol{f},\boldsymbol{b},\boldsymbol{p}),$ (1)
where the estimated image is reconstructed by $\boldsymbol{\hat{x}}=\mathcal{D}(\boldsymbol{\hat{z}})$ with the corresponding image decoder.
Forward Process. The noised feature maps are
$\boldsymbol{z_{t_{z}}}=\alpha_{t_{z}}\mathbf{z}+\sigma_{t_{z}}\boldsymbol{\epsilon}_{t_{z}},~{}~{}~{}\boldsymbol{m_{t_{m}}}=\alpha_{t_{m}}\mathbf{m}+\sigma_{t_{m}}\boldsymbol{\epsilon}_{t_{m}},$ (2)
where $\epsilon_{z}$ and $\epsilon_{m}$ $\sim\mathbf{N}(0,I)$ are three independently sampled Gaussian noise, and the time step $t_{z}$ and $t_{m}$ are sampled from a uniform distribution $\mathcal{U}[0,1]$.
Reverse Process. The denoiser $\hat{\boldsymbol{\epsilon}}_{\theta}$ simultaneously denoises both $\boldsymbol{z_{t_{z}}}$ and $\boldsymbol{m_{t_{m}}}$, which can be trained with:
$\mathcal{L}=\mathbb{E}\left[\|\boldsymbol{\epsilon}_{z}-\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{z}_{t_{z}};\mathbf{f},\mathbf{b},\mathbf{p}))\|^{2}\right]+\mathbb{E}\left[\|\boldsymbol{\epsilon}_{m}-\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{m_{t_{m}}};\mathbf{f},\mathbf{b},\mathbf{p})\|^{2}\right].$
The denoiser model consists of the UNet model and the three encoders for the foreground object $\boldsymbol{f}$, the background scene $\boldsymbol{b}$, and the position prompt $\boldsymbol{p}$, respectively.
Figure 2: Mask-aware Dual Diffusion Model (MADD). The RGB image feature $\mathbf{z}$ and object mask $\mathbf{m}$ are jointly denoised, conditioning on the embeddings of the foreground object $\mathbf{f}$, background object $\mathbf{b}$, and the prompt $\mathbf{p}$. (green: reverse process $t\rightarrow t\mathrm{-}1$)
4.2 Architecture Details
UNet Model. We utilize a single UNet with two expert input-output branches to denoise them simultaneously. The two tasks share the entire UNet except for the conv_in, first Down Block, last Up Block, and the conv_out. These independent blocks serve as the Expert input-output branch. Skip connection is also performed between the corresponding Expert input and output branches.
Foreground Encoder. In MADD, foreground images serve as a guidance condition and are incorporated into the model via cross-attention. The foreground embedding is extracted using a ViT-like visual encoder. While SD employs the CLIP [40] encoder for text-aligned semantics, we follow Efficient-3DiM [20] and use DINOv2 [37] to preserve fine-grained object details, making it more suitable for our task.
Background Encoder. Following SD-based image editing methods [44, 5], we employ a pre-trained VAE encoder to convert the background image into a 4-channel latent representation. Since the background is pixel-aligned with the output, we concatenate its latent map with $\mathbf{z_{t}}$ before feeding them into the U-Net. The pre-trained VAE decoder then reconstructs the final RGB image from the latent output.
Position Prompt Encoder. We design a unified representation to process different position prompts. Sparse and dense prompts are converted into a 1-channel position map, guiding foreground object placement. Point prompts are transformed into Gaussian heatmaps, bounding boxes are filled with ones inside and zeros outside, and masks are binary maps. For null prompts, an all-one image implies all positions are possible. The position map is resized to match the latent map’s spatial dimensions and concatenated with it before being fed into the diffusion U-Net.
4.3 Implementation Details
We initialize our model with a pre-trained Stable Diffusion Inpainting model and apply strong data augmentation to enhance affordance learning. To optimize fine-tuning, we use a coarse-to-fine strategy: training at $128\times 128$ resolution (batch size 1024) for 35K steps, then fine-tuning at $256\times 256$ (batch size 256) for 15K steps.
5 Experiments
5.1 Results on the SAM-FB Test Set
[…]
mask bbox mask bbox mask bbox
 [42] 15.41 15.47 0.7079 0.8058 860 883
 [49] 33.68 24.59 0.7664 2373 1615
 [30] - 13.82 0.7896 817
 [24] 14.49 14.42 0.8014 0.8637 857 845
Ours 13.53 13.60 0.8727 0.8658 760 775
Evaluation Metrics. Following the previous work [35, 19, 22, 6], we employ the FID score to measure the quality of images obtained by generative models, MSE for pixel similarity and the CLIP score to evaluate the semantic similarity between the edited region and the reference foreground.
Quantitative Results. We compare our method with Stable Diffusion [42], PBE [49], GLI-GEN [30], and Human Affordance [24] on the SAM-FB test set to show the effectiveness of our method. For fairness, we fine-tuned GLI-GEN [28] on SAM-FB training data for 2k steps with captions generated by ViT-GPT2 [39, 9] and use Box + Text + Image Inpainting mode to adapt it. Since Human Affordance [24] does not release training data or model weights, we re-trained their model on SAM-FB. The results in Table 3 show that our method attains the best FID score and the highest CLIP score, illustrating our model can be a simple yet strong baseline for affordance-aware object insertion task. Table 4 shows that the mask prompt achieves the best results as it provides more accurate position information.
Prompt Mask Bbox Point Null Avg.
FID 13.53 13.60 13.66 13.96 13.69
MSE 760 775 772 860 792
CLIP Score 0.8727 0.8658 0.8567 0.8034 0.8497
Table 4: Comparison of position prompts on the SAM-FB test set.
Figure 4: We test ambiguous prompts (points and blank) on in-the-wild images. With point prompts, our model adjusts foreground properties for affordance-aware insertion, while it autonomously finds suitable positions when no prompt is given.
Qualitative Results. We left 3786 images as test split. Figure 3 presents the visualization results on the SAM-FB test set. In each group, the leftmost image depicts the background marked with a position prompt. Our MADD predicts the RGB image and mask of the inserted object, which are shown in the last two images of each group. These results demonstrate that MADD not only inserts objects with high quality but also accurately predicts object masks.
Prompt Non-Null Null
IoU ($\uparrow$) 0.5405 0.4696
Table 5: Mask evaluation
Model Components
Method FID ($\downarrow$) CLIP ($\uparrow$)
Baseline 18.32 0.79
+ Dual diffusion 14.43 0.84
+ Expert branch 13.69 0.85
Higher Resolution
256 $\times$ 256 13.69 0.85
512 $\times$ 512 10.66 0.82
Table 6: Ablation study results on SAM-FB test set.
Mask Evaluation. Table 5 shows the IoU of the mask between the predicted mask and the ground truth on the SAM-FB test set. The relatively high IoU shows that our method learns effective affordance information.
Computational complexity. We access the computational complexity for the newly added modules including Dual Diffusion and Expert Branch. Using Human Affordance model with DINOv2 as baseline, it has 863.02M parameters with 28.3B FLOPs. Our Expert branches shared most of the parameters with the original baseline model and end up with 873.75M parameters with 31.1B FLOPs.
Figure 5: MADD can work on images of higher resolution, generating sharper edges, clearer reflections, improved texture details.
5.2 Ablation Studies
Model Components. Table 6 presents the results of our ablation study. For model design, we replace the image encoder in the Human Affordance model with DINOv2 as the baseline. Next, we diffuse RGB image and object mask simultaneously, but sharing the entire UNet. Finally, we use two Expert branches for mask and RGB streams. The results show that all of them improved performance. Metrics are averaged across different position prompts. Please refer to the supplementary for results on each prompt type.
Higher Resolution. Even though we trained our model at a resolution of 256x256, it is also capable of handling higher-resolution inputs, such as 512x512. To prove that, we then fine-tuned our model on SAM-FB at a resolution of 512x512 for only 200 steps and compared it with the original checkpoint on the test split. Figure 5 shows the results of the original checkpoint and the fine-tuned version at higher resolution using the same input. We highlighted the details of the inserted object. It is clearly demonstrated that the model generates sharper edges, clearer reflections, and improved texture details after fine-tuning at higher resolution. The CLIP score drops slightly at 512×512, likely due to the increased complexity of higher-resolution foregrounds, making fine-grained detail synthesis more challenging.
5.3 Results on In-the-wild Images
We evaluate our model on in-the-wild web-crawled images. Ambiguous Prompts. Training with position augmentations enables the model to capture affordance relationships between objects and scenes. MADD refines object position, size, and view for coherence with the background.
In Figure 4, MADD adjusts a person’s position to stand on the ground rather than floating, aligns a car’s orientation with the lane, and scales coffee beans to match the scene. The model can also place objects without explicit prompts and generate diverse, reasonable insertions based on user input and learned affordance, as shown in Figure 6.
Human Evaluation. To test the generalization ability of our MADD model, we performed affordance insertion on in-the-wild images and compared the results with Stable Diffusion XL [38], GLI-GEN [30], PBE [49] and ObjectStitch [44]. Instead of merely relying on metrics like FID and CLIP Score, we also conducted user study to achieve more comprehensive comparison. We asked 10 users to rank 10 groups of composited image generated from from different model according to the following criteria: 1) Foreground and Background Integration; 2) Foreground Clarity and Detail; 3) Foreground Appearance Consistency with Reference; 4) Lighting and shade on Foreground; 5) Color Consistency. Figure 7 shows the distribution of rank for different models, where rank 1 and rank 5 represent the best and worst quality respectively. Our model achieves 50% of Rank-1 place and 1.60% of the Rank-5 place, which outperforms other methods. Please refer to supplementary for details of each criteria. Our model consistently achieved the highest proportion of Rank-1 placements across all evaluation criteria, as indicated by the largest segments in each pie chart. Especially for keeping the consistency of foreground appearance with reference image. This dominance in Rank-1 distribution across multiple criteria highlights the model’s superior performance compared to others in all aspects.
Figure 7: Rank distribution for different methods. Our method has the most proportion of rank 1 and least proportion of rank 5.
Figure 8: More in-the-wild affordance-insertion examples. The model can generate an affordance-feasible solution to insert the foreground objects according to the background scene.
More Results. Figure 8 presents additional in-the-wild examples of affordance-aware object insertion with various position prompts. For ambiguous prompts, such as points or null inputs, MADD infers a reasonable position and adjusts the foreground accordingly. For bounding boxes and masks, MADD treats them as guidance rather than strict constraints, refining placements to ensure coherence. Please refer to the supplementary for additional examples that further demonstrate MADD’s ability to generate suitable and diverse insertions in real-world scenarios.
6 Conclusion
In this paper, we extend the concept of affordance beyond human-centered tasks by introducing affordance-aware object insertion, enabling seamless object placement into scenes while adhering to affordance principles. To support this, we construct SAM-FB, a dataset of 3.16 million foreground-background pairs spanning 3,000+ object categories. We propose MADD, a dual-stream diffusion model that jointly denoises foreground appearance and object masks, leveraging affordance relationships for context-aware insertion. MADD accommodates various position prompts and can infer placements autonomously. Our model generalizes well across diverse objects and outperforms prior diffusion-based composition methods, achieving coherent and affordance-aware insertions.
Supplementary Material
Appendix A More Implementation Details
A.1 SAM-FB Dataset
Figure 9(a) shows the object masks before and after the data quality control. We see that with our designed data quality control, the foreground object masks have better quality. Figure 9(b) illustrates the word cloud of our SAM-FB dataset, we observe that our SAM-FB dataset consists of diverse object categories.
A.2 Traning Details
We constructed our dual-stream UNet based on the Stable Diffusion Inpainting v1.5 model, incorporating several modifications. Specifically, we replicated both the first down-sampling block (including conv_in and the first Down Block) and the last up-sampling block (including the last Up Block and conv_out) to accommodate dual-stream inputs and outputs. These independent blocks serve as the expertise input-output branch. Skip connection is also performed between the corresponding expertise input and output branches. To bypass the need to start from scratch, we initialized our model using the pre-trained Stable Diffusion Inpainting v1.5 checkpoint for the unchanged blocks. Fine-tuning is resource-intensive. To optimize our computing resources, we employed a gradual scaling-up approach to train our model. Initially, the model was trained at a resolution of $128\times 128$, with a batch size of 1024 and a learning rate of $1.25\times 10^{-4}$ for 35K steps on 2 A-100 GPUs. This phase included 5,000 warm-up steps, followed by a constant schedule. We then fine-tuned the model at a higher resolution of $256\times 256$, reducing the batch size to 256 and adjusting the learning rate to $5\times 10^{-5}$. For the diffusion process, the time step is 1000 with a linear noise scheduler. We use 2048 samples of SAM-FB for testing and the rest for training. For each foreground object, all different position prompts $\boldsymbol{p}$ can be generated with the mask $\boldsymbol{s}$. We apply classifier-free guidance [17], dropping all conditions with 0.1 probability. To ensure a fair comparison with other methods, we re-implemented and re-trained the closest method, Human Affordance, since the authors did not release either the model weights or the full dataset they used. We adopted the diffusion model architecture provided by the authors and trained it using our SAM-FB dataset. Additionally, we replaced the mask input in the original model with our position map to ensure compatibility with the SAM-FB dataset.
A.3 Data Augmentation
@@
Figure 9: 9(a) shows the candidate foreground samples in the pipeline. The upper row shows four low-quality samples. The lower row shows the samples after data quality control. 9(b) shows the word cloud of foreground categories in the SAM-FB dataset.
@@
Baseline 16.05 17.31 18.73 21.18 18.32 0.8209 0.8254 0.7768 0.7509 0.7935
+Dual Diffusion 13.91 14.12 14.58 14.92 14.38 0.8582 0.8578 0.8435 0.8032 0.8407
+Expertise branch 13.53 13.60 13.66 13.96 13.69 0.8727 0.8658 0.8567 0.8034 0.8497
@@ B.1 Evaluation metrics
Figure 11: Rank-1 distribution for each criterion. Each pie chart represents the proportion of times each model achieved Rank-1 for a specific evaluation criterion. Our method dominates every metric.
@@ B.2 Human Evaluation
We conducted our user survey by asking 10 users to rank outputs from different model in 10 affordance insertion settings, therefore gaining 100 data points.
Figure 12: Affordance insertion away from inpainted region
B.3 Effect of Inpaintings
Figure 12 shows results if the inserted location is away from masked region. It illustrates that the effects of inpainting artifacts are limited.
@@ C.1 Details Maintaining
Using DINOv2 features, the model preserves the appearance more effectively than other image editing models, especially for objects with detailed textures. Figure 10 compares our model with other image editing models. Our model demonstrates the ability to retain the details of an object’s appearance, even texture on the object.
@@ C.2 SAM-FB Test Split
Figure 13 shows a visualization result with different baselines on SAM-FB test split. Our model generated more authentic results compared to other baseline models.
Figure 13: Samples on SAM-FB test split. Our model inserted the bag with authentic appearance.
@@ C.3 Detailed Ablations
Figure 14: Example of in-the-wild insertion results with details. Our model could keep the appearance better and adjust the foreground’s properties better compared with SD [42], GLI-GEN [30] and PBE [49] on common objects. In the last row, the model generated reasonable insertion when provided ambiguous prompts.
Training Datasets FID ($\downarrow$) CLIP ($\uparrow$)
COCO 31.17 0.75
SAM-FB 28.45 0.77
Table 8: Ablation study results on SAM-FB test set.
C.4 Training Datasets
To illustrate SAM-FB helps the model on affordance insertion task, we retrained our model on 50k COCO2014 and SAM-FB subset, each for 10k steps, then evaluated on PascalVOC as in-the-wild benchmark. Table 8 shows SAM-FB lead to lower FID and higher CLIP score.
C.5 In-the-wild Generalization
We compare our methods with other baseline models on in-the-wild images, and Figure 14 shows the visualization results for comparison. In the first five rows, we performed affordance insertion task providing a bounding box on both common objects like apple and uncommon objects like cruta. Our methods generated the images with highest quality, authentic to the reference foreground with proper lighting. In the last row, we show more affordance insertion results when providing ambiguous prompts. The model will automatically find the proper affordance relationship and adjust the location and view of the foreground object. It’s notable that in the first case, when inserting the cruise, it is actually the view from the back of the reference image.
Method Anydoor Ours
CLIP Score ($\uparrow$) 0.8209 0.8658
DINO Score ($\uparrow$) 0.667 0.693
RfHE($\downarrow$) 2.18 2.21
Table 9: Anydoor comparison
Figure 15: Affordance insertion compare with Anydoor
C.6 Comparison with Anydoor
Table 9 reports quantitative results on SAM-FB and Rank from Human Evaluation (RfHE) on in-the-wild images from web (Unsplash). Users prefer MADD for more reasonable insertion (Figure 15). Anydoor introduces artifacts around objects.
C.7 Video Demo
Please refer to the video demo in the attachment.
C.8 Failure Cases
Figure 16 shows some failure cases when using the model to perform affordance insertion. Generally, when prompted null position, it requires the model to search for a possible position to insert the object. However, if there are already similar objects in the scene e.g., traffic light in the first row, it is easy to mislead the model and end up with inserting nothing. When the background is too complex and the foreground object is too small, such as a vegetable in a supermarket shown in the second row, it is also difficult for the model to insert the object correctly.